a salesperson has two job offers. Company A offers a weekly salary of $630 plus commission of 14% of sales.Company B offers a weekly salary of $1260 plus commission of 7% of sales. What is the amount of sales above which Company A's offer is the better of the two?
Answer by rfer(16322)

630+0.14x=1260+0.07x
0.07x=1260-630
0.07x=630
x=630/0.07
x=$9000